OpenStudy (anonymous):

Simplify: a(x + y) - a(x - y) Question 3 options: 2ax 2ay ax+ay 2ax-2ay

OpenStudy (anonymous):

im really bad need help!

OpenStudy (freckles):

here you don't have like terms I suggest distributing then combining like terms

OpenStudy (freckles):

and what I mean by you don't have like terms is that x+y is not the same as x-y both terms have different variable parts as is

OpenStudy (freckles):

use the rule I mentioned on your earlier post

OpenStudy (freckles):

a(b+c)=ab+ac

OpenStudy (anonymous):

freckles?

OpenStudy (anonymous):

will u please give me the answer.. D: i will fan and medal

OpenStudy (anonymous):

@Vocaloid please help me

Vocaloid (vocaloid):

hint: A(B+C) = A*B + A*C a(x+y) = ????

OpenStudy (calculusxy):

\[a(x+y) - a(x-y) \implies ax +ay -ax +ay\] combine the like terms

OpenStudy (calculusxy):

like terms are the variables that are raised to the same power

oregonduck (oregonduck):

after combining like terms you get d 2ax-2ay

OpenStudy (calculusxy):

\[ax + ay - ax + ay\] \[(ax - ax) \implies ? \] \[(ay + ay) \implies ? \]

OpenStudy (calculusxy):

did you find your answer?

OpenStudy (anonymous):

is the answer c @calculusxy